karen mixed together 9 gal of brand A fruit drink and 3 gal of brand B fruit drink which contains 16% fruit juice find the percent of fruit in Brand A if the mixture contained 25% fruit juice
1
Expert's answer
2012-09-27T10:39:42-0400
Let brand A contains X L of juice. Then

X + 3*0.16 = (9 + 3)*0.25

and so, X = 2.52. Therefore, the part of the fruit juice in brand A is

P = 2.52/9 = 0.28

and brand A fruit drink contains 28% of juice.